# Can Large Language Models Act as Ensembler for Multi-GNNs?

### v2

@@ Appendix B Baselines
Table 9: Detailed hyperparameters of aligning multi-GNNs.
Model Dataset Hidden size GNN representation size Number of layers Learning rate of GNN encoder Initial learning rate of the classifier Dropout Rate Weight Decay
@@
Model Dataset Lora dropout Loraplus lr ratio Training batch size Learning rate Early stop epoch Warmup ratio
@@
Appendix C Experiment setup
Appendix D Prompt design
The prompt is divided into two parts: Instruction and Input. The former specifies what the LLM should do with the input, defines the format of the input, and highlights the characteristics of the input content. The latter provides the specific text of target node or graph and their multiple GNN representations. An example of a prompt template can be seen in Figure 5. For graph classification, we follow MolXPT Liu et al. (2023) to design prompts for molecular graphs as shown in Figure 6.
Appendix E Detailed hyperparameter settings
The details on the setting of hyperparameters in the first and second stages of LensGNN  are shown in Tables 9 and 10, respectively. The hyperparameters for the 20-shot experiment are shown in Tables 11 and 12.
Model Dataset Hidden size GNN representation size Number of layers Learning rate of GNN encoder Initial learning rate of the classifier Dropout Rate Weight Decay
@@ Appendix E Detailed hyperparameter settings
Model Dataset Lora dropout Loraplus lr ratio Training batch size Learning rate Early stop epoch Warmup ratio

### v3

@@ Appendix C Experiment setup
Table 9: Hyperparameter Search Spaces.
Appendix D Prompt design
The prompt design follows three key principles:
  1. 1.
  2. 2.
  3. 3.
The prompt is divided into two parts: Instruction and Input. The former specifies what the LLM should do with the input, defines the format of the input, and highlights the characteristics of the input content. The latter provides the specific text of target node or graph and their multiple GNN representations.
An example of the prompt template can be seen in Figure 5. For graph classification, we follow MolXPT Liu et al. (2023) to design prompts for molecular graphs as shown in Figure 6.
Appendix E Best variants
The best variants in Table 2 and Table 3 are listed in Table 10 and Table 11, respectively.
Backbone LLM Cora PubMed ogbn-arXiv
InternLM2.5-7B LensGNN-ALL LensGNN-ALL LensGNN-[GAT+GIN]
Baichuan2-7B LensGNN-[GCN+GAT] LensGNN-ALL LensGNN-ALL
Baichuan2-13B LensGNN-[GCN+GIN] LensGNN-ALL LensGNN-ALL
Table 10: Best variants in Table 2.
Backbone LLM Cora PubMed
Baichuan2-7B LensGNN-ALL LensGNN-ALL
InternLM2.5-7B LensGNN-ALL LensGNN-[GAT+GIN]
Table 11: Best variants in Table 3.
Appendix F One-shot results
We further conduct experiments on one-shot setting, the results are presented in Table 12. From the table, we see that
Model Cora PubMed Citeseer Wiki-CS
GCN 0.1890 0.4155 0.3242 0.2873
GAT 0.2543 0.4993 0.3506 0.4094
GIN 0.2021 0.4071 0.2763 0.3278
LensGNN 0.5938 0.8379 0.5332 0.3047
Table 12: Classification accuracy in one-shot setting.
our method still achieves outstanding performance on the Cora, PubMed, and Citeseer datasets, which shows the strong generalizability. For WikiCS, due to the very scare labeled data, GNN ensembling is adversely affected.
Appendix G Traditional methods ensembling LLM and GNNs
We have conducted additional experiments employing traditional ensemble learning methods to ensemble the predictions from both the LLM and GNNs (GCN, GAT and GIN) on 4 datasets. The results are as Table 13:
Method LLM Cora PubMed Citeseer Wiki-CS
Bagging Baichuan2-13B 86.90 86.16 75.55 76.47
Stacking Baichuan2-13B 85.24 85.70 70.53 67.35
LensGNN* Baichuan2-13B 91.88 95.68 79.31 83.47
Table 13: Comparison with traditional ensemble of LLM and GNNs. (%)
As demonstrated in our experiments, LensGNN consistently outperforms all conventional ensemble methods.
Appendix H LLama2-7B as backbone
We have supplemented our experiments by implementing LLaMA2-7B as the backbone LLM. We conducted these experiments on a medium-scale dataset, PubMed, as a representative, with results presented in Table 14:
Model PubMed
MLP 82.40
GCN 83.78
GAT 83.59
GIN 82.03
GraphSAGE 87.01
Graphormer 88.75
BERT 93.91
SentenceBERT 92.49
DeBERTa 93.45
GraphGPT-MIX-7B 64.76
LLaGA-HO-7B(GENERAL) 75.01
InstructGLM-Llama-7B 93.84
DGTL 87.10
SNS-GPT4 93.80
GAugLLM 74.15
Baichuan2-13B 36.04
LensGNN*-LLaMA2-7B 94.59
Table 14: Performance comparison (%) on PubMed dataset using LLaMA2-7B as backbone. (%)
Experimental results demonstrate that our model with LLaMA2-7B as the backbone also achieves superior accuracy compared to: (1) GNN baselines, (2) LM-based methods, and (3) other LLM-based approaches.
Appendix I Ensemble more GNN
We further incorporate a larger number of GNNs to evaluate how LensGNN scales and whether the performance continues to improve. The results are presented in Table 15. The experiment is based on the backbone of Internlm2.5-7B-chat, with the number of graph tokens $t=4$. Overall, we see that with the increase in the number of GNNs, the model performance improves further. However, from the results on Cora and Pubmed, simply adding GNNs will not bring sustained improvement. We speculate that this is because more GNNs will introduce more graph tokens and also noise. The increased input length will bring difficulty in model learning, so does noise.
Based on the empirical findings presented in Table 15, we propose a heuristic recommendation for researchers to select 3-4 high-performing GNNs (e.g., GCN, GAT, GIN) when constructing ensembles for new datasets. Notably, the inclusion of additional GNN models incurs minimal computational overhead, allowing for flexible expansion until validation accuracy approaches saturation—the point at which further performance improvements become negligible. This strategy balances model diversity and efficiency while maximizing predictive accuracy.
Variants of LensGNN Number of GNNs Cora PubMed Citeseer Wiki-CS
LensGNN-GCN 1 0.9059 0.9490 0.7836 0.8294
LensGNN-[GCN+GAT] 2 0.9003 0.9503 0.7868 0.8306
LensGNN-[GCN+GAT+GIN] 3 0.9114 0.9508 0.7884 0.8334
LensGNN-[GCN+GAT+GIN+GraphSAGE] 4 0.9095 0.9513 0.7852 0.8332
LensGNN-[GCN+GAT+GIN+APPNP] 4 0.9040 0.9500 0.7915 0.8366
LensGNN-[GCN+GAT+GIN+GraphSAGE+APPNP] 5 0.9022 0.9472 0.7931 0.8417
Table 15: Classification accuracy when integrating more GNNs.
Appendix J Additional hyperparameter sensitivity analysis
We further conduct hyperparameter sensitivity analysis on Citeseer and Wiki-CS, whose results are given in Figure 7. From the figure, we see that the model can also achieve good results at $t=8$ in terms of the overall performance, which is similar as the results in Figure 4.
Figure 7: Hyperparameter sensitivity analysis.
Appendix K Input length analysis
We investigate the impact of LLM input length on model efficiency and performance across Cora, Citeseer, and Wiki-CS. Specifically, we established different limits on total number of input tokens. When exceeded, some neighbor text is randomly truncated. Within this constraint, more text introduces additional noise but also provides richer useful information, leading to better model performance—albeit at increased computational cost. The experimental results are presented in Figure 8.
Figure 8: Input length analysis.
Our experiments using Baichuan-7B as the backbone record sample statistics with total token length limits ranging from 0.5K to 1.75K, including average input string length, training time, and test accuracy. While Cora’s shorter texts show negligible variations, both Citeseer and Wiki-CS exhibit a clear trend: longer inputs increased training time but also improved accuracy.
Appendix L Detailed hyperparameter settings
The details on the setting of hyperparameters in the first and second stages of LensGNN are shown in Tables 16 and 17, respectively. The hyperparameters for the 20-shot experiment are shown in Tables 18 and 19.
Model Dataset Hidden size
GNN representation
size
Number of
layers
Learning rate of
GNN encoder
Initial learning rate of
the classifier
Dropout Rate Weight Decay
@@ Appendix I Ensemble more GNN
Table 16: Detailed hyperparameters of aligning multi-GNNs.
Model Dataset
Lora dropout
Loraplus
lr ratio
Training
batch size
Learning rate
Early stop
epoch
Warmup ratio
@@
Model Dataset Hidden size
GNN representation
size
Number of
layers
Learning rate of
GNN encoder
Initial learning rate of
the classifier
Dropout Rate Weight Decay
@@
Model Dataset
Lora dropout
Loraplus
lr ratio
Training
batch size
Learning rate
Early stop
epoch
Warmup ratio